simpify:
(-x^3-x+9) + (x+2)

thanks

Found 2 solutions by orca, edjones:
Answer by orca(409)

First remove the parenthesis, then combine like terms.
(-x^3-x+9) + (x+2)
= -x^3 - x + 9 + x + 2
= -x^3 -x + x +9 + 2
= -x^3 + 11

Answer by edjones(8007)


(-x^3-x+9) + (x+2)
=-x^3+11
.
Ed